OpenStudy (anonymous):

1 over 4 (2x − 14) = 4

OpenStudy (anonymous):



OpenStudy (anonymous):

cross multiply and find x good luck

OpenStudy (whpalmer4):

(2x-14) is in the denominator? \[\frac{1}{4(2x-14)} = 4\]? Not \[\frac{1}{4}(2x-14) = 4\]?

OpenStudy (whpalmer4):

@BSANDSTORM which one is it?